PFI-ZEKE PHOTOELECTRON SPECTROSCOPY OF ALKALI METAL-ETHER COMPLEXES

\underline{BRADFORD R. SOHNLEIN}, JASON F. FULLER, SHENGGANG LI, and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40508.

This talk will present the PFI-ZEKE photoelectron spectra of two alkali metal-ether complexes: Li(DME) and Li(DXE); DME = CH_{3}OCH_{3}, DXE = CH_{3}O(CH_{2})_{2}OCH_{3}. Both complexes were produced in a pulsed molecular beam and identified by time-of-flight mass spectrometry. The ZEKE spectra of these complexes show adiabatic ionization potentials and vibrational structures consisting of metal-ligand and ligand-based vibrations. The spectroscopic measurements were combined with density functional theory and Franck-Condon factor calculations to establish ground electronic states and geometries of the neutral and ionic species. For Li(DME), the neutral and ionic complexes were determined to be in C_{2v} symmetry with Li binding to oxygen. For Li(DXE), the Li atom binds to both oxygen atoms in a C_{2} five-membered ring structure, with the O-C-C-O dihedral angle of ∼50^{o} in both the ion and neutral species. The ground electronic states of the DME complex and its ion are ^{2}A_{1} and ^{1}A_{1}, whereas those of the corresponding DXE complex and its ion are ^{2}A and 1A, respectively. Furthermore, the binding energies of the two neutral complexes were derived using a thermodynamic cycle. This simple calculation requires the ionization energies of the metal atom and complex as well as the dissociation energy of the cation complex.